OpenStudy (anonymous):

Using substitution y = 3x + 4 y = -x + 8

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

its right! thank you so much

OpenStudy (anonymous):

y=3x+4............(1) y=-x+8............(2) substitute the value of y from 2nd equation in 1st:: -x+8=3x+4 4=4x x=1 therefore ,y=3(1)+4 y=7

OpenStudy (anonymous):

Thanks @nitz i needed that too!

OpenStudy (anonymous):

welcome.......